A typical car driven 14,000 miles a year emits to the atmosphere about 12 kg per year of NOx (nitrogen oxides), which causes smog in major population areas. Natural gas burned in the furnace emits about 4.6 g of NOx per therm, and the electric power plants emit about 7.2 g of NOx per kWh of electricity produced. Consider a household that has 2 cars and consumes 10,000 kWh of electricity and 1,400 therms of natural gas. Determine the amount of NOx emission to the atmosphere per year for which this household is responsible.
Answer #1

NOx emission from cars:

There are two cars, one car produces 12kg per year, therefore two cars produced 2 × 12 = 24 kg per year.

NOx emission from burning natural gas:

One therm produces 4.6g of NOx, then 1400 therms produced 1400 × 4.6g = 6440g = 6.44 kg of NOx.

NOx emission for 1kWh is 7.2g, then for 10,000kWh is 10,000 × 7.2g = 72000g = 72 kg.

Amount of NOx emission in a year for which this household is responsible is 24kg + 6.44kg + 72kg = 102.44kg